### v1

# Robust Preference Optimization: Aligning Language Models with Noisy Preference Feedback

Abstract
Standard human preference-based alignment methods, such as Reinforcement Learning from Human Feedback (RLHF), are a cornerstone technology for aligning Large Language Models (LLMs) with human values. However, these methods are all underpinned by a critical, yet flawed assumption: human preferences are homogeneous (representing a single, unified preference) and the collected data is noiseless (free from error). In reality, neither is true since human preference is pluralistic and annotators can make mistakes. This creates a discrepancy between the recorded data and the ground-truth preferences, which can misguide the model and degrade its performance. To address this challenge, we introduce Robust Preference Optimization (RPO). RPO employs an Expectation-Maximization (EM) algorithm to infer the posterior probability of each label’s correctness, which is used to adaptively re-weigh each data point in the training loss to mitigate noise. We further generalize this approach by establishing a theoretical link between arbitrary preference losses and their corresponding probabilistic models. This generalization enables the systematic transformation of existing alignment algorithms into their robust counterparts, elevating RPO from a specific algorithm to a meta-framework for robust preference alignment. Theoretically, we prove that under the condition of a perfectly calibrated model, RPO is guaranteed to converge to the true noise level of the dataset. Our experiments demonstrate RPO’s effectiveness as a meta-framework, consistently enhancing four state-of-the-art alignment algorithms (DPO, IPO, SimPO, and CPO). When applied to Mistral and Llama 3 models, the RPO-enhanced methods achieve substantial win rate gains on AlpacaEval 2 and Arena-Hard, with improvements of up to 7.0% and 5.4%, respectively.
1 Introduction
Aligning Large Language Models (LLMs) with human values is a critical prerequisite for developing safe and reliable AI systems. Reinforcement Learning from Human Feedback (RLHF) has emerged as the dominant paradigm for this task (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). To mitigate the complexity and instability of the traditional RLHF pipeline, simpler and more direct methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) have been developed, which reframe alignment as a classification-like problem.
Figure 1: The Robust Preference Optimization (RPO) framework iteratively learns a collective preference from noisy data using an Expectation-Maximization (EM) algorithm. In each cycle, the E-step computes a confidence score for each label by inferring the posterior probability of the true underlying preference. The M-step then uses these scores as adaptive weights to update both the LLM policy and the annotator reliability parameters. This process repeats, allowing the policy and reliability scores to mutually improve.
However, these alignment methods implicitly assume that preference datasets provide a faithful representation of a single, ground-truth set of human values. This assumption is often violated in practice for two primary reasons. The first is simple annotator error, stemming from inattention or misunderstanding. The second, more subtle source is the inherent pluralism of human preferences (Berlin, 2013; Bakker et al., 2022). For many subjective topics, legitimate disagreements arise from diverse annotator perspectives. From the perspective of an algorithm seeking to learn a single, unified preference model, these structured disagreements are functionally equivalent to noise. The impact of such noise is severe, studies indicate that a substantial portion (20% - 40%) of preference pairs in alignment datasets can be noisy (Gao et al., 2024; Natarajan et al., 2013). An algorithm that naively trusts every label with equal weight is therefore highly susceptible to being misled by incorrect feedback. A mere 10 percentage point increase in the noise rate can lead to a 30 percentage point drop in a model’s win rate (Gao et al., 2024).
To address this challenge, we propose Robust Preference Optimization (RPO). Instead of treating labels as a fixed truth, our approach aims to learn a robust model of collective preference that best explains the entire noisy and diverse dataset. The core innovation of RPO is its departure from the hard labels used in traditional RLHF. We treat the true preference as a latent variable and compute soft preference labels to estimate the underlying consensus. To achieve this, RPO employs an Expectation-Maximization (EM) framework that simultaneously models annotator reliability while optimizing the LLM. In the E-step, it infers the posterior probability that each annotated label is correct, effectively estimating annotator reliability. In the M-step, it uses these probabilities as adaptive weights to update the LLM, thereby learning from a dynamically re-weighted preference signal.
Our comprehensive experiments validate RPO as a highly effective and versatile meta-framework. We show that applying RPO consistently enhances four state-of-the-art alignment algorithms (DPO, IPO, SimPO, and CPO) across two different base models (Mistral-7B and Llama-3-8B) on two widely-used benchmarks (Table 2). RPO-enhanced methods achieve substantial win-rate gains over their standard counterparts, with improvements of up to 7.0% on AlpacaEval 2 and 5.4% on Arena-Hard. Furthermore, we theoretically prove that RPO can recover the true reliability of annotators (Theorem 4.2) and empirically verify this guarantee in controlled experiments (section 5.4).
In summary, our contributions are as follows:
  • We propose Robust Preference Optimization, an algorithm that achieves robust LLM alignment by actively denoising preference data. RPO formulates this as a latent variable problem, employing an EM approach to infer the reliability of each label while simultaneously optimizing the LLM policy.
  • We theoretically establish a generalized RPO framework by using the Gibbs distribution to connect any preference loss function to an underlying probabilistic model. This generalization elevates RPO to a meta-framework, capable of systematically transforming existing alignment methods into their robust counterparts.
  • We conduct extensive experiments that demonstrate the practical effectiveness and versatility of our framework. By applying RPO to several popular alignment methods, we show significant performance improvements, providing a unified approach to robust LLM alignment.
2 Related work
LLM alignment with hard preference labels.
The standard paradigm for aligning Large Language models (LLMs) with human values is Reinforcement Learning from Human Feedback (RLHF), which involves training a reward model and then fine-tuning the policy against it (Christiano et al., 2017; Ouyang et al., 2022). To mitigate the complexity and instability of this multi-stage process, a family of simpler, direct alignment algorithms has emerged (Rafailov et al., 2023; Azar et al., 2023; Meng et al., 2024; Hong et al., 2024). These methods bypass the explicit reward modeling stage by optimizing a direct classification-style loss on the preference data. However, a critical limitation shared by these methods is their reliance on hard preference labels. This approach models human feedback as a definitive, binary choice, treating every label with equal and absolute confidence. Consequently, it is highly vulnerable to the significant label noise present in real-world datasets, as standard loss functions can lead models to overfit to corrupted labels (Natarajan et al., 2013; Zhang and Sabuncu, 2018). A simple annotation error, such as an accidental misclick, is given the same weight as a deliberate, high-quality judgment. This inability to distinguish between reliable feedback and noise means that the model’s performance degrades significantly as the error rate increases (Frénay and Verleysen, 2013; Gao et al., 2024). In contrast, a soft preference label approach, which models preferences probabilistically, can better handle this uncertainty. By assigning a confidence score or weight to each label, the model can learn to rely more on high-quality signals while down-weighting probable noise (Müller et al., 2019; Song et al., 2024). This is the core advantage of our work; RPO outperforms prior methods because its soft-label framework explicitly models and mitigates the impact of label noise, leading to a more robust and effective alignment.
Learning from noisy feedback.
The vulnerability to label noise situates preference alignment within the classic machine learning problem of Learning with Noisy Labels (LNL) (Natarajan et al., 2013). Foundational work in LNL, such as the Dawid and Skene model (Dawid and Skene, 1979), used an EM algorithm to simultaneously infer true labels while estimating annotator reliability. This principle was later extended to pairwise comparisons in the Crowd-BT model (Chen et al., 2013), which jointly estimates item scores and annotator reliability in crowdsourced ranking tasks. In the context of modern LLM alignment, several approaches have recently been proposed to address label noise. These can be broadly categorized into methods that design robust loss functions and those that implement data-centric filtering strategies. In the first category, rDPO (Chowdhury et al., 2024) constructs an unbiased estimator of the true loss but requires the global noise rate to be known a priori. Hölder-DPO (Fujisawa et al., 2025) introduces a loss with a ”redescending property,” which inherently nullifies the influence of extreme outliers without needing a known noise rate. In the second category, Selective DPO (Gao et al., 2025) proposes filtering examples based on their difficulty relative to the model’s capacity—a concept orthogonal to label correctness—using validation loss as a proxy. Another recent framework, ORPO (Hong et al., 2024), employs a hybrid iterative approach that alternates between training with a robust loss and explicitly filtering out a fixed fraction of high-loss samples. Our proposed RPO framework is distinct from these contemporary methods. Instead of merely mitigating the consequences of noise (e.g., high-loss samples), RPO addresses its root cause by directly modeling the data-generating process. By treating annotator reliability as a latent variable to be inferred, it can distinguish signal from noise in a principled manner, offering a more robust path to alignment.
3 Methodology
Table 1: Formulations of the preference loss ($\mathcal{L}_{\text{pref}}$) for prominent alignment algorithms.
Method Preference Loss $\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l})$
DPO (Rafailov et al., 2023) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
IPO (Azar et al., 2023) $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\beta}\right)^{2}$
SimPO (Meng et al., 2024) $\log\sigma(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{w}|x)-\gamma)$
CPO (Xu et al., 2024) $\log\sigma(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{w}|x))-\log\pi_{\theta}(y_{w}|x)$
This section details our proposed RPO algorithm. We begin by briefly reviewing the standard DPO framework, which serves as our baseline in section 3.1. We then identify its key limitation in handling noisy data, which motivates the core of our work: a new framework that treats preferences as latent variables. Finally, we derive the RPO algorithm using an Expectation-Maximization (EM) approach and discuss its practical implementation.
3.1 Preliminaries: Direct Preference Optimization
The goal of preference alignment is to fine-tune a language model policy, $\pi_{\theta}$, using a dataset of preferences $\mathcal{D}=\{(x,y_{w},y_{l})_{i}\}_{i=1}^{N}$, where response $y_{w}$ is preferred over $y_{l}$ for a given prompt $x$. Direct Preference Optimization (DPO) (Rafailov et al., 2023) offers a simple and effective method for this, bypassing the complex multi-stage pipeline of traditional RLHF (Christiano et al., 2017; Ouyang et al., 2022). DPO directly optimizes the policy by minimizing a simple classification loss:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
where $\sigma(\cdot)$ is the sigmoid function, $\pi_{\text{ref}}$ is a fixed reference policy and $\beta$ is a scaling hyperparameter.
3.2 RPO Framework: Core Assumptions
A critical limitation of DPO is its implicit assumption that all observed preferences in $\mathcal{D}$ are correct. In practice, this data is often noisy. To address this, we propose Robust Preference Optimization (RPO), which is built upon two core assumptions that reframe the problem.
Assumption 1: A latent collective preference.
The cornerstone of our framework is the assumption of a latent collective preference, denoted $y_{w}\succ^{\ast}y_{l}$. This represents the idealized collective consensus we aim for the LLM to learn. The observed preferences in the dataset, $y_{w}\succ_{k}y_{l}$ (from annotator $k$), are merely noisy signals of this unobservable ground truth. To bridge this gap, we introduce a binary latent variable $z_{i}\in\{0,1\}$ for each data point, where $z_{i}=1$ if the annotator’s label correctly aligns with the collective preference, and $z_{i}=0$ otherwise. This allows us to formally model the reliability of each annotator $k$ with a parameter $\eta_{k}\triangleq p(z_{i}=1|k_{i}=k)$. Here, $k_{i}$represents the index of the annotator for the $i$-th data point.
Input: Dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i},k_{i})\}_{i=1}^{N}$; Base policy $\pi_{\theta}$, reference policy $\pi_{\text{ref}}$; Preference loss $\mathcal{L}_{\text{pref}}$; Hyperparameters: learning rate $\lambda$, epochs $E$, EMA momentum $\alpha$, initial annotator reliabilities $\eta_{k}\in[0.5,1]$ for all $k\in\{1,\dots,K\}$
1
2for epoch = 11 to $E$ do
3 for batch $\mathcal{B}\subset\mathcal{D}$ do
    // // E-Step: Calculate label confidences for the current batch
4      For each sample $i\in\mathcal{B}$, compute $w_{i}$ using current $\theta$ and $\eta_{k_{i}}$ via (4);
5
    // // M-Step: Update parameters based on the batch
6
7     Compute the weighted loss $\mathcal{L}_{\text{RPO}}(\theta)$ for the batch via (5);
8      Update parameters $\theta$ using an optimizer (e.g., AdamW (Loshchilov and Hutter, 2019));
9
10    for each annotator $k$ present in the batch do
11         Update $\eta_{k}$ via (7);
12
13      end for
14
15   end for
16
17 end for
Algorithm 1 Robust Preference Optimization (RPO)
Assumption 2: A general probabilistic model for preferences.
Building on this latent variable model, we must also define the probability of the ground-truth preference itself, $p(y_{w}\succ^{\ast}y_{l}|x,\theta)$. To accommodate various preference losses beyond DPO (e.g., IPO (Azar et al., 2023)), our framework is designed to work with any preference loss function, $\mathcal{L}_{\text{pref}}$. Table 1 provides several examples of such loss functions used in prominent alignment algorithms.
To connect these diverse loss functions to a unified probabilistic interpretation, we draw inspiration from the Boltzmann distribution (Luce, 1959). We assume that for any preference loss function $\mathcal{L}_{\text{pref}}$, the probability of a preference is proportional to the exponentiated negative loss $\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))$. This yields a general definition for the ground-truth preference probability:
$p(y_{w}\succ^{\ast}y_{l}|x,\theta)=\sigma\left(\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w};\theta)-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l};\theta)\right),$ (2)
where $\sigma(\cdot)$ is the sigmoid function. This formulation converts any preference loss into a well-defined probability distribution. For instance, with the standard DPO loss, this equation recovers the Bradley-Terry model (Bradley and Terry, 1952) (see Appendices A and B for derivations).
3.3 The RPO Algorithm via Expectation-Maximization
Based on these core assumptions, we aim to find the parameters $\theta$ and $\boldsymbol{\eta}$ that maximize the marginal log-likelihood of the observed data. The probability of a single observed preference is obtained by marginalizing over the latent variable $z_{i}$:
$p(y_{w,i}\succ_{k_{i}}y_{l,i}|x_{i},\theta,\boldsymbol{\eta})=p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)\eta_{k_{i}}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)(1-\eta_{k_{i}}).$ (3)
Directly maximizing $\sum_{i}\log p(y_{w,i}\succ_{k_{i}}y_{l,i})$ is intractable due to the sum inside the logarithm. We therefore employ the Expectation-Maximization (EM) algorithm (see detailed derivation in Appendix C), which iterates between two steps. In this iterative process, the superscript $(t)$ will denote the values of parameters at iteration $t$.
E-Step: Inferring label correctness.
In the E-step, given the current parameters $\theta^{(t)}$ and $\boldsymbol{\eta}^{(t)}$, we compute the posterior probability $w_{i}$ that the $i$-th observed label is correct. This value $w_{i}$ acts as a ”soft label” or the model’s confidence in the data point.
$w_{i}^{(t)}\leftarrow\frac{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}}{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})(1-\eta_{k_{i}}^{(t)})}.$ (4)
where $p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})$ and $p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})$ can be computed according to (2).
M-Step: weighted parameter update.
In the M-step, we update the policy parameters $\theta$ and reliabilities $\boldsymbol{\eta}$ using the confidences $w_{i}^{(t)}$ computed in the E-step. This step conveniently separates into two independent updates.
First, the policy is updated by minimizing a weighted loss function. As established in Assumption 2, our probabilistic model for $p(y_{w}\succ^{\ast}y_{l})$ allows RPO to work with any preference loss $\mathcal{L}_{\text{pref}}$, making it a versatile meta-framework. The general RPO loss is:
$\min_{\theta}\mathcal{L}_{\text{RPO}}(\theta)=-\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)\right]$ (5)
Second, the reliability $\eta_{k}$ for each annotator is updated to the average confidence of all labels they provided. This has a simple and efficient closed-form solution:
$\eta_{k}^{(t+1)}\leftarrow\frac{\sum_{i\in\mathcal{I}_{k}}w_{i}^{(t)}}{N_{k}}$ (6)
Here we define the index set of labeled pairs as $\mathcal{I}_{k}=\{\,i:k_{i}=k\,\}$, and the number of labels as $N_{k}$.
3.4 Practical implementation with mini-batch training
While the exact M-step updates are clear, performing a full iteration over the entire dataset to re-calculate the annotator reliabilities $\boldsymbol{\eta}$ after each policy update step can be computationally expensive. To balance computational efficiency and performance, we introduce a more practical online update for $\eta_{k}$ using an Exponential Moving Average (EMA). Instead of a hard assignment, we perform a soft update based on the statistics from the current mini-batch $\mathcal{B}$:
$\eta_{k}\leftarrow(1-\alpha)\eta_{k}+\alpha\cdot\frac{\sum_{i\in\mathcal{B}\cap\mathcal{I}_{k}}w_{i}}{N_{k,\mathcal{B}}}$ (7)
Here, $N_{k,\mathcal{B}}$ is the number of examples from annotator $k$ in the current mini-batch, and $\alpha\in(0,1]$ is a momentum hyperparameter. The complete training procedure for RPO is summarized in Algorithm 1:
4 Theoretical analysis of RPO
The robustness of RPO stems from its adaptive weighting mechanism. This section first provides an intuitive analysis of these training dynamics and then formalizes this intuition with theoretical guarantees, demonstrating that the RPO framework can recover the true reliability of annotators.
At the start of training, when the language model is not yet well-optimized, its predictions are uncertain, and the probabilities $p(y_{w}\succ^{\ast}y_{l}|x,\theta)$ are close to 0.5. The confidence score $w_{i}$ approximates the annotator’s reliability, $\eta_{k_{i}}$. The loss then acts as a form of label smoothing, preventing the model from being severely misled by incorrect labels early on. As the policy improves, its behavior adapts. For a high-quality label, the model predicts a high probability for the winning response, and $w_{i}$ approaches 1, causing the loss to function like a standard preference optimization objective. Conversely, $w_{i}$ approaches 0 for a noisy label. The loss is then dominated by the $(1-w_{i})$ term, which flips the optimization direction toward the true preference.
We now formalize the intuition that RPO can recover the true reliability of annotators. We provide this analysis under an idealized setting: full-batch training where the M-step for the policy parameters $\theta$ is assumed to have converged perfectly. While our practical implementation in Algorithm 1 uses mini-batch gradient updates (a form of Generalized EM), this idealized analysis provides a strong theoretical justification for our framework.
We first show that under a well-specified preference model, the EM update for $\eta_{k}$ takes the form of an averaging operator $T_{k}(\eta)$. The following lemma identifies the true reliability as its fixed point.
Lemma 4.1 (Fixed point of the EM operator).
Let $\theta^{\star}$ be a perfectly calibrated parameter such that the model distribution matches the ground-truth preference distribution: $p(y_{w,i}\succ^{*}y_{l,i}|x_{i})=p(y_{w,i}\succ^{\ast}y_{l,i}\mid x_{i},\theta^{\star}).$ Denote $p_{i}^{*}=p(y_{w,i}\succ^{*}y_{l,i}|x_{i})$, the EM update for annotator $k$ takes the form
$T_{k}(\eta)\triangleq\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}\frac{p_{i}^{\star}\eta}{p_{i}^{\star}\eta+(1-p_{i}^{\star})(1-\eta)}.$
Then the true reliability $\eta_{k}^{*}\triangleq\mathbb{E}[z_{i}\mid k_{i}=k]$ is a fixed point of this operator, i.e. $T_{k}(\eta_{k}^{*})=\eta_{k}^{*}.$
The proof is shown in Appendix D. This lemma confirms the consistency of our formulation, showing that the true parameter $\eta_{k}^{*}$ is a stationary point under the EM dynamics. This naturally raises the question of whether this fixed point is unique and attractive, which the following theorem addresses. We next show that this fixed point is unique and globally attractive for the EM iterations.
Theorem 4.2 (Identification and convergence of RPO).
Under the calibration and observation model of Lemma 4.1, assume that not all $p_{i}^{\star}$ equal $\tfrac{1}{2}$ for $i\in\mathcal{I}_{k}$. Consider the full-batch EM iterations $\eta_{k}^{(t+1)}=T_{k}(\eta_{k}^{(t)})$ with $T_{k}$ as in Lemma 4.1. Then for any initialization $\eta_{k}^{(0)}\in(0,1)$ the EM iterates converge to the true reliability $\eta_{k}^{*}$: $\lim_{t\to\infty}\eta_{k}^{(t)}=\eta_{k}^{\star}.$
The proof is provided in Appendix E. Although Algorithm 1 employs mini-batch updates, Theorem 4.2 analyzes the full-batch idealization and should be viewed as a mean-field guide; in section 5.4 we empirically corroborate that the mini-batch procedure closely tracks this behavior and recovers accurate reliability estimates in practice.
5 Experiments
Table 2: Performance comparison on AlpacaEval 2 (Dubois et al., 2024) and Arena-Hard (Li et al., 2024). LC and WR denote length-controlled and raw win rate, respectively. Each pair consists of a standard alignment algorithm and its robust counterpart using our RPO framework (e.g., R-DPO). We also report the performance of the base models without any fine-tuning for reference. Our RPO-enhanced methods consistently outperform the base algorithms. The best scores in each pairing are in bold, with absolute gains from RPO highlighted in red.
Method Mistral-Instruct (7B) Llama-3-Instruct (8B)
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC (%) WR (%) WR (%) LC (%) WR (%) WR (%)
Base Model 21.1 16.5 10.0 29.7 29.9 19.5
DPO 28.5 28.6 12.4 40.8 42.9 23.4
R-DPO (Ours) 35.5 (+7.0) 33.0 (+4.4) 14.7 (+2.3) 44.1 (+3.3) 46.2 (+3.3) 28.8 (+5.4)
IPO 30.8 28.0 11.5 43.6 41.6 27.0
R-IPO (Ours) 32.9 (+2.1) 30.5 (+2.5) 12.0 (+0.5) 48.3 (+3.3) 48.6 (+7.0) 29.0 (+2.0)
SimPO 28.3 29.7 13.6 44.5 37.1 20.8
R-SimPO (Ours) 30.4 (+2.1) 32.9 (+3.2) 15.6 (+2.0) 46.9 (+2.4) 39.4 (+2.3) 23.1 (+2.5)
CPO 26.3 26.4 13.5 35.9 40.3 25.7
R-CPO (Ours) 27.6 (+1.3) 27.8 (+1.4) 14.7 (+1.2) 40.1 (+4.2) 43.8 (+3.5) 28.3 (+2.6)
In this section, we conduct a comprehensive set of experiments to evaluate the preformance of RPO. We begin in section 5.1 by detailing our experimental setup, including the models, datasets, evaluation benchmarks, and baseline algorithms. In Section 5.2, we present our main results. We then conduct an ablation study in Section 5.3 to analyze the framework’s sensitivity to its key hyperparameters. In section 5.4, we provide an empirical verification of our theoretical claims from Theorem 4.2.
5.1 Experimental setup
Models and training settings.
We use two state-of-the-art open-source large language models as our base models: Mistral-7B-Instruct-v0.2 and Meta-Llama-3-8B-Instruct. For fine-tuning, we utilize two datasets from the SimPO paper (Meng et al., 2024), which were generated via on-policy sampling using prompts from the UltraFeedback dataset (Cui et al., 2024). The specific datasets are mistral-instruct-ultrafeedback for the Mistral model and llama3-ultrafeedback-armorm for the Llama-3 model.  [^1] See Appendix F for links to models and datasets. As these datasets do not provide annotator-specific information, we model the preferences as if they originate from a single, virtual annotator (K=1K=1). [^2] This is a reasonable simplification. For instance, a pool of two annotators with reliabilities $\eta_{A}$ and $\eta_{B}$, appearing with frequencies $p_{A}$ and $p_{B}$ respectively, can be modeled as a single annotator with an equivalent reliability $\eta_{\text{unified}}=p_{A}\eta_{A}+p_{B}\eta_{B}$.
Evaluation benchmarks.
We assess model performance on two widely recognized evaluation benchmarks. The first is AlpacaEval 2 (Dubois et al., 2024), an automatic, LLM-based evaluator that measures model performance by computing the win rate against reference outputs. It provides both a raw Win Rate (WR) and a Length-Controlled (LC) Win Rate to account for verbosity bias. The second is Arena-Hard (Li et al., 2024), a challenging benchmark composed of difficult prompts crowdsourced from the LMSYS Chatbot Arena. It is designed to differentiate high-performing models by testing them on complex, real-world user queries. Performance is reported as the win rate against a suite of other models.
Table 3: Ablation study on the initial annotator reliability ($\eta_{0}$) and the EMA momentum ($\alpha$). The best-performing settings used in our main experiments are highlighted.
Metric Initial $\boldsymbol{\eta_{0}}$ EMA $\boldsymbol{\alpha}$
0.99 0.9 (Ours) 0.75 0.55 0.001 0.01 0.1 (Ours) 0.5 1.0
AlpacaEval2 LC (%) 30.9 35.5 31.1 31.4 30.9 30.1 35.5 33.4 31.1
AlpacaEval2 WR (%) 31.7 33.0 33.3 32.0 27.8 27.2 33.0 34.8 28.9
Arena-Hard WR (%) 12.3 14.7 12.4 11.8 12.9 13.6 14.7 14.0 12.8
Implementation details.
All experiments were conducted using the Hugging Face TRL library (von Werra et al., 2020). We performed full-parameter fine-tuning for all models, leveraging DeepSpeed ZeRO-3 (Rasley et al., 2020) to manage memory across our hardware setup. Models were trained for a single epoch using the AdamW optimizer (Loshchilov and Hutter, 2019) with a cosine learning rate scheduler, a peak learning rate of $5\times 10^{-7}$, and a global batch size of 64. We also maintained an Exponential Moving Average (EMA) of the model parameters with a decay rate of $\alpha=0.1$. The temperature for DPO and its variants was set to $\beta=0.01$. For all RPO variants, the reliability parameter was initialized to $\eta^{(0)}=0.9$ and was updated via the E-step after the training epoch was complete. All training was performed on a cluster of 8x NVIDIA A800 40GB GPUs.
Baseline algorithms.
To demonstrate that RPO operates as a versatile meta-framework, we benchmark it against four popular direct preference alignment methods: DPO (Rafailov et al., 2023); IPO (Azar et al., 2023), which uses a squared hinge loss to optimize preferences; SimPO (Meng et al., 2024), which proposes a simplified, reference-free reward formulation normalized by sequence length; and CPO (Xu et al., 2024), which adds a term to directly maximize the likelihood of the preferred response. For each of these baselines, whose loss functions are detailed in Table 1, we compare the original algorithm to its RPO-enhanced counterpart (e.g., DPO vs. R-DPO).
5.2 Main results
As shown in Table 2, our experimental results provide strong evidence for the effectiveness of the RPO framework. We summarize the key findings below.
RPO as a meta-framework.
The primary takeaway from our experiments is that RPO functions as a highly effective and generalizable meta-framework. Across all four baseline algorithms (DPO, IPO, SimPO, CPO), both base models (Mistral-7B and Llama-3-8B), and all evaluation benchmarks, the RPO-enhanced version consistently and significantly outperforms its standard counterpart. For instance, R-DPO improves upon DPO by up to +5.4 points on Arena-Hard, and R-IPO surpasses IPO by +7.0 points in win rate on AlpacaEval 2. This universal performance lift validates that RPO is a robust wrapper that can be applied to a wide range of existing preference alignment methods to enhance their resilience to label noise.
Some algorithms benefit more than others.
While all methods benefit from RPO, the magnitude of the improvement varies, shedding light on the interplay between different loss functions and our robust training scheme. RPO delivers particularly substantial gains to DPO and IPO, especially on the more capable Llama-3 model. This suggests that algorithms with standard loss formulations (like DPO’s cross-entropy) may be more vulnerable to label noise and thus benefit more from RPO’s adaptive re-weighting.
Scaling effects with base LLM.
An interesting insight is that the benefits of RPO appear to scale with the capability of the base model. The performance gains are consistently more pronounced on the Llama-3-8B model than on Mistral-7B. For example, the average improvement on the challenging Arena-Hard benchmark is more than double on Llama-3 compared to Mistral. More powerful models are better equipped to capitalize on the cleaner training signal that RPO provides by identifying and down-weighting noisy data.
5.3 Ablation study
We conduct an ablation study to analyze the sensitivity of RPO to two key hyperparameters: the initial annotator reliability, $\eta_{0}$, and the EMA momentum parameter, $\alpha$. All experiments are performed using the R-DPO algorithm on the Mistral-7B-Instruct-v0.2 model. The results are summarized in Table 3.
(a) Single-annotator.
(b) Two-annotator.
Figure 2: Empirical verification of annotator reliability estimation. Ground-truth reliability ($\eta$ GPT-4o) is established using GPT-4o’s labels. Different reliability levels are simulated by injecting synthetic noise into copies of the UltraFeedback dataset. In the two-annotator setting (b), Annotator 1 is the baseline with no added noise, while noise is progressively added to Annotator 2’s data. The plots confirm that the RPO-estimated reliability ($\eta$ RPO, dashed lines) closely tracks the ground-truth values (solid lines).
Effect of initial $\eta_{0}$.
The initial reliability $\eta_{0}$ sets the model’s prior belief about the correctness of the labels in the dataset. As shown in Table 3, the model’s performance is best when $\eta_{0}$ is set to 0.9, which was the value used in our main experiments. An overly optimistic initialization (e.g., $\eta_{0}=0.99$) can cause the model to trust noisy labels too strongly at the beginning of training, hindering the denoising process. Conversely, a pessimistic initialization (e.g., $\eta_{0}=0.55$) treats the data as highly unreliable from the outset, which can slow down the model’s ability to learn the underlying collective preference. An initial value of 0.9 appears to strike the right balance, starting with a reasonable assumption of data quality.
Effect of EMA parameter $\alpha$.
The EMA parameter $\alpha$ governs the update rate of the annotator reliability scores, balancing the influence of historical estimates against new information from the current mini-batch. Our experiments confirm that the optimal performance is achieved with $\alpha=0.1$. The model shows considerable sensitivity to this parameter. A very small $\alpha$ (e.g., 0.001) makes the reliability updates exceedingly slow, preventing the estimates from adapting to the model’s evolving understanding of the data. On the other hand, a very large $\alpha$ (e.g., 1.0) makes the updates highly volatile, as the reliability score becomes dependent solely on the samples in the current mini-batch. This can lead to unstable training dynamics. A balanced value of 0.1 ensures that the reliability estimates are updated smoothly and stably.
5.4 Empirical verification of Theorem 4.2
We conduct controlled experiments to verify Theorem 4.2. Our setup is designed to align with the theorem’s assumption of a perfectly calibrated model, for which we use a small-scale base model, Qwen2.5-0.5B-Instruct  [^3] See Appendix F for links to models., to ensure fast convergence. To simulate annotators with varying levels of reliability, we create distinct copies of the UltraFeedback dataset (Cui et al., 2024) for each annotator and inject a controlled degree of synthetic noise into their respective dataset.
We test two scenarios, with results presented in Figure 2: (a) Single Annotator: A single annotator whose dataset is modified with a synthetically controlled noise rate. (b) Two Annotators: A scenario with two annotators, where Annotator 1 serves as a baseline using the original data without added noise, while the dataset for Annotator 2 is injected with progressively increasing noise levels.
The results in Figure 2 show that the estimated reliability $\eta$ (RPO) closely tracks the ground-truth $\eta$ (GPT-4o) in both single-annotator (Figure 2(a)) and two-annotator (Figure 2(b)) settings. Notably, in the two-annotator experiment, RPO successfully identifies the stable reliability of the baseline annotator while accurately tracking the declining reliability of the noisy one.
6 Conclusion
In this paper, we introduce Robust Preference Optimization (RPO), a novel framework designed to address the critical challenge of aligning LLMs with noisy human preference data. Our approach is distinct from existing methods as it employs an Expectation-Maximization algorithm to infer the reliability of each preference pair, treating labels as soft, dynamic weights rather than fixed ground truths. As a meta-framework, RPO consistently enhances multiple state-of-the-art alignment algorithms, achieving significant performance gains (up to a 7.0% win rate increase on AlpacaEval 2) across various base models.
Acknowledgments and Disclosure of Funding
This research was supported by National Natural Science Foundation of China (No.62325405,62406159), Tsinghua University Initiative Scientific Research Program, Tsinghua-Efort Joint Research Center for EAI Computation and Perception, Beijing National Research Center for Information Science, Technology (BNRist), Beijing Innovation Center for Future Chips, and State Key laboratory of Space Network and Communications.
Appendix A Derivation of general probabilistic model
Here we provide the detailed derivation for (2). For a given prompt $x$ and candidate responses $y_{w},y_{l}$, we assume the probability of the ground-truth preference $y_{w}\succ^{\ast}y_{l}$ is proportional to $\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))$. That is:
$p(y_{w}\succ^{\ast}y_{l}|x,\theta)\propto\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))$ (8)
Similarly, for the inverse preference:
$p(y_{l}\succ^{\ast}y_{w}|x,\theta)\propto\exp(-\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w}))$ (9)
Since $y_{w}\succ^{\ast}y_{l}$ and $y_{l}\succ^{\ast}y_{w}$ are the only two mutually exclusive outcomes for a binary preference, their probabilities must sum to 1. Using the property of normalized probabilities from a proportional relationship, we have:
$p(y_{w}\succ^{\ast}y_{l}|x,\theta)$ $=\frac{\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))}{\exp(-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l}))+\exp(-\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w}))}$
$=\frac{1}{1+\exp(-(\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w})-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l})))}$
$=\sigma\left(\mathcal{L}_{\text{pref}}(x,y_{l}\succ y_{w})-\mathcal{L}_{\text{pref}}(x,y_{w}\succ y_{l})\right)$
The last line is the General Probabilistic Model in (2).
Appendix B Consistency with Bradley-Terry model for DPO
We show that (2) is consistent with the Bradley-Terry model when applied to DPO. The DPO loss for a preferred pair $(y_{w},y_{l})$ given prompt $x$ is:
$\mathcal{L}_{\text{DPO}}(x,y_{w}\succ y_{l})=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ (10)
Let $S(x,y_{w},y_{l})=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$. Then, we can write:
$\mathcal{L}_{\text{DPO}}(x,y_{w}\succ y_{l})$ $=-\log\sigma(S(x,y_{w},y_{l}))$
$\mathcal{L}_{\text{DPO}}(x,y_{l}\succ y_{w})$ $=-\log\sigma(S(x,y_{l},y_{w}))=-\log\sigma(-S(x,y_{w},y_{l}))$
Substituting these into our general probabilistic model ((2)):
$p(y_{w}\succ^{\ast}y_{l}|x,\theta)$ $=\sigma\left(\mathcal{L}_{\text{DPO}}(x,y_{l}\succ y_{w})-\mathcal{L}_{\text{DPO}}(x,y_{w}\succ y_{l})\right)$
$=\sigma\left(\log\sigma(S(x,y_{w},y_{l}))-\log\sigma(-S(x,y_{w},y_{l}))\right)$
$=\sigma\left(\log\frac{\sigma(S(x,y_{w},y_{l}))}{1-\sigma(S(x,y_{w},y_{l}))}\right)$
$=\sigma(S(x,y_{w},y_{l}))$
$=\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
This resulting probability exactly matches the form of the Bradley-Terry model [Bradley and Terry, 1952] for preferences, where the implicit reward of a response $y$ is $r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$.
Appendix C Derivation of the RPO EM algorithm
The primary objective of Robust Preference Optimization (RPO) is to find the model parameters $\theta$ and the vector of annotator reliabilities $\boldsymbol{\eta}$ that maximize the log-likelihood of the observed data. The observed data consists of prompts, chosen and rejected responses, and the annotator’s index, denoted as $X=\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i},k_{i})\}_{i=1}^{N}$.
The log-likelihood function is given by:
$\mathcal{L}(\theta,\boldsymbol{\eta})=\sum_{i=1}^{N}\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)\eta_{k_{i}}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)(1-\eta_{k_{i}})\right]$ (11)
There is a sum inside the logarithm, which makes direct optimization intractable. The Expectation-Maximization (EM) algorithm is an iterative procedure designed to solve such maximum likelihood problems with latent variables by alternating between an Expectation (E) step and a Maximization (M) step.
C.1 Derivation of the Q-Function (The E-Step)
The EM algorithm simplifies the problem by working with the complete data, $(X,Z)$, where $Z=\{z_{i}\}_{i=1}^{N}$ is the set of all latent variables.
The complete-data log-likelihood, $\mathcal{L}_{c}$, assumes that we know the values of all latent variables $z_{i}$:
$\mathcal{L}_{c}(\theta,\boldsymbol{\eta};X,Z)=\sum_{i=1}^{N}\left(z_{i}\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)\eta_{k_{i}}\right]+(1-z_{i})\log\left[p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)(1-\eta_{k_{i}})\right]\right)$ (12)
This form is tractable because the logarithm acts on products, which can be separated into sums.
The core idea of EM is to iteratively maximize the expectation of the complete-data log-likelihood. This expectation, known as the Q-function, is taken with respect to the posterior distribution of the latent variables $Z$, given the observed data $X$ and the parameter estimates from the current iteration, $(\theta^{(t)},\boldsymbol{\eta}^{(t)})$.
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})\equiv\mathbb{E}_{Z|X,\theta^{(t)},\boldsymbol{\eta}^{(t)}}[\mathcal{L}_{c}(\theta,\boldsymbol{\eta};X,Z)]$ (13)
To compute this expectation, we push the expectation operator inside the summation. The only random variables in $\mathcal{L}_{c}$ are the $z_{i}$.
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})=\sum_{i=1}^{N}\left(\mathbb{E}[z_{i}]\log\left[p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)\eta_{k_{i}}\right]+(1-\mathbb{E}[z_{i}])\log\left[p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)(1-\eta_{k_{i}})\right]\right)$ (14)
The term $\mathbb{E}[z_{i}]$ is the expectation of the binary variable $z_{i}$, which is its posterior probability of being 1. This probability is conditioned on the observed data and the parameters from the current iteration $t$. We denote this posterior probability as $w_{i}^{(t)}$, which is computed in the E-Step:
$w_{i}^{(t)}$ $\equiv\mathbb{E}[z_{i}|X_{i},\theta^{(t)},\boldsymbol{\eta}^{(t)}]$
$=p(z_{i}=1|y_{w,i}\succ_{k_{i}}y_{l,i},x_{i},\theta^{(t)},\boldsymbol{\eta}^{(t)})$
$=\frac{p(y_{w,i}\succ_{k_{i}}y_{l,i}|z_{i}=1,x_{i},\theta^{(t)})p(z_{i}=1|k_{i},\boldsymbol{\eta}^{(t)})}{p(y_{w,i}\succ_{k_{i}}y_{l,i}|x_{i},\theta^{(t)},\boldsymbol{\eta}^{(t)})}$
$=\frac{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}}{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})(1-\eta_{k_{i}}^{(t)})}$ (15)
Substituting $w_{i}^{(t)}$ into the expression yields the final form of the Q-function:
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})=\sum_{i=1}^{N}\left[w_{i}^{(t)}\log(p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)\eta_{k_{i}})+(1-w_{i}^{(t)})\log(p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)(1-\eta_{k_{i}}))\right]$ (16)
C.2 Deriving the RPO Framework (The M-Step)
The goal of the M-Step is to find the parameters for the next iteration, $(\theta^{(t+1)},\boldsymbol{\eta}^{(t+1)})$, by maximizing the Q-function that was constructed using the parameters from the current iteration $t$.
$(\theta^{(t+1)},\boldsymbol{\eta}^{(t+1)})=\arg\max_{\theta,\boldsymbol{\eta}}Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})$ (17)
To perform this maximization, we can first expand the Q-function by separating the terms involving the policy $\theta$ from those involving the annotator reliabilities $\boldsymbol{\eta}$.
$Q(\theta,\boldsymbol{\eta}|\theta^{(t)},\boldsymbol{\eta}^{(t)})=$ $\underbrace{\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)\right]}_{\text{Depends only on }\theta}$
$+\underbrace{\sum_{i=1}^{N}\left[w_{i}^{(t)}\log\eta_{k_{i}}+(1-w_{i}^{(t)})\log(1-\eta_{k_{i}})\right]}_{\text{Depends only on }\boldsymbol{\eta}}$ (18)
Because the Q-function is separable into two independent parts, we can maximize each part separately to find the new parameters.
To find the optimal $\theta^{(t+1)}$, we hold $\boldsymbol{\eta}$ fixed and maximize the terms in the Q-function that depend on $\theta$:
$\theta^{(t+1)}$ $=\arg\max_{\theta}\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)\right]$
$=\arg\min_{\theta}\left(-\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|\theta)\right]\right)$ (19)
The expression inside the $\arg\min$ is precisely the weighted RPO loss function, $\mathcal{L}_{\text{RPO}}(\theta)$. This establishes that the M-step for the policy parameters is equivalent to minimizing this weighted loss, using weights $w_{i}^{(t)}$ from the E-step.
To find the optimal $\eta_{k}^{(t+1)}$ for a specific annotator $k$, we hold $\theta$ fixed and maximize the terms in the Q-function relevant to $\eta_{k}$. These terms only involve samples labeled by annotator $k$ (where $k_{i}=k$):
$\eta_{k}^{(t+1)}=\arg\max_{\eta_{k}\in[0,1]}\sum_{i:k_{i}=k}\left[w_{i}^{(t)}\log\eta_{k}+(1-w_{i}^{(t)})\log(1-\eta_{k})\right]$ (20)
To find the maximum, we take the derivative with respect to $\eta_{k}$ and set it to zero:
$\frac{\partial}{\partial\eta_{k}}\sum_{i:k_{i}=k}\left[w_{i}^{(t)}\log\eta_{k}+(1-w_{i}^{(t)})\log(1-\eta_{k})\right]$ $=0$ (21)
$\sum_{i:k_{i}=k}\left[\frac{w_{i}^{(t)}}{\eta_{k}}-\frac{1-w_{i}^{(t)}}{1-\eta_{k}}\right]$ $=0$ (22)
$\frac{1}{\eta_{k}}\sum_{i:k_{i}=k}w_{i}^{(t)}$ $=\frac{1}{1-\eta_{k}}\sum_{i:k_{i}=k}(1-w_{i}^{(t)})$ (23)
$\frac{1}{\eta_{k}}\sum_{i:k_{i}=k}w_{i}^{(t)}$ $=\frac{1}{1-\eta_{k}}\left(N_{k}-\sum_{i:k_{i}=k}w_{i}^{(t)}\right)$ (24)
where $N_{k}$ is the total number of annotations provided by annotator $k$. Cross-multiplying gives:
$(1-\eta_{k})\sum_{i:k_{i}=k}w_{i}^{(t)}$ $=\eta_{k}\left(N_{k}-\sum_{i:k_{i}=k}w_{i}^{(t)}\right)$ (25)
$\sum_{i:k_{i}=k}w_{i}^{(t)}-\eta_{k}\sum_{i:k_{i}=k}w_{i}^{(t)}$ $=\eta_{k}N_{k}-\eta_{k}\sum_{i:k_{i}=k}w_{i}^{(t)}$ (26)
$\sum_{i:k_{i}=k}w_{i}^{(t)}$ $=\eta_{k}N_{k}$ (27)
This yields the intuitive and closed-form update rule for the reliability at iteration $t+1$:
$\eta_{k}^{(t+1)}=\frac{\sum_{i:k_{i}=k}w_{i}^{(t)}}{N_{k}}$ (28)
This shows that the updated reliability for an annotator is simply the average posterior probability (or confidence) from the previous iteration that their labels were correct.
Appendix D Proof of Lemma 4.1
Fix an annotator $k$ and write $p_{i}^{\star}=p(y_{w,i}\succ^{\ast}y_{l,i}\mid x_{i})$; by perfect calibration, $p(y_{w,i}\succ^{\ast}y_{l,i}\mid x_{i},\theta^{\star})=p_{i}^{\star}$ and $p(y_{l,i}\succ^{\ast}y_{w,i}\mid x_{i},\theta^{\star})=1-p_{i}^{\star}$. Under the generative model, the observed label for pair $i$ arises from a mixture over $z_{i}\in\{0,1\}$ with mixing weights $\eta$ and $1-\eta$, hence Bayes’ rule gives the E-step posterior
$w_{i}(\eta)=p\!\left(z_{i}=1\mid y_{w,i}\succ^{k}y_{l,i},\,x_{i},\,\theta^{\star},\,\eta\right)=\frac{p_{i}^{\star}\,\eta}{p_{i}^{\star}\,\eta+(1-p_{i}^{\star})(1-\eta)}.$
By EM, $T_{k}(\eta)=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}w_{i}(\eta)$, yielding the stated form of $T_{k}$. Now set $\eta=\eta_{k}^{\star}=\mathbb{E}[z_{i}\mid k_{i}=k]$. Using the same Bayes expression with $\eta$ replaced by $\eta_{k}^{\star}$ and taking conditional expectation over the observed sample for fixed $i$,
$w_{i}(\eta_{k}^{\star})=p\!\left(z_{i}=1\mid y_{w,i}\succ^{k}y_{l,i},\,x_{i},\,\theta^{\star},\,\eta_{k}^{\star}\right)=\mathbb{E}\!\left[z_{i}\mid y_{w,i}\succ^{k}y_{l,i},\,x_{i},\,\theta^{\star}\right].$
Averaging over $i\in\mathcal{I}_{k}$ and applying the law of total expectation,
$T_{k}(\eta_{k}^{\star})=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}w_{i}(\eta_{k}^{\star})=\mathbb{E}[z_{i}\mid k_{i}=k]=\eta_{k}^{\star},$
so $\eta_{k}^{\star}$ is a fixed point of $T_{k}$.
Appendix E Proof of Theorem 4.2
We use the same notation as in Appendix D. To begin, we define the function $\ell_{k}(\eta)$. The term $\ell_{k}(\eta)$ represents the observed-data log-likelihood for the set of labels provided by annotator $k$, under the idealized assumption of a perfectly calibrated model. It is formally defined as:
$\ell_{k}(\eta)=\sum_{i\in\mathcal{I}_{k}}\log\left(p_{i}^{\star}\eta+(1-p_{i}^{\star})(1-\eta)\right).$
The significance of this function is that it serves as the objective that the full-batch EM algorithm seeks to maximize. The remainder of this proof demonstrates that the EM updates guarantee a monotonic increase in $\ell_{k}(\eta)$ and that its unique maximum corresponds to the true annotator reliability, $\eta_{k}^{\star}$.
For $\eta\in(0,1)$, we can write the denominator of the posterior probability as $d_{i}(\eta)=(1-p_{i}^{\star})+(2p_{i}^{\star}-1)\eta=(1-\eta)(1-p_{i}^{\star})+\eta\,p_{i}^{\star}>0$. Differentiating $\ell_{k}(\eta)$ with respect to $\eta$ yields:
$\ell_{k}^{\prime}(\eta)=\sum_{i\in\mathcal{I}_{k}}\frac{2p_{i}^{\star}-1}{d_{i}(\eta)},\qquad\ell_{k}^{\prime\prime}(\eta)=-\sum_{i\in\mathcal{I}_{k}}\frac{(2p_{i}^{\star}-1)^{2}}{d_{i}(\eta)^{2}}.$
If at least one $i$ satisfies $p_{i}^{\star}\neq\tfrac{1}{2}$, then one summand in $\ell_{k}^{\prime\prime}$ is strictly negative on $(0,1)$ and the rest are $\leq 0$, hence $\ell_{k}^{\prime\prime}(\eta)<0$ on $(0,1)$. Thus, $\ell_{k}$ is strictly concave and admits a unique stationary point $\widehat{\eta}_{k}$, which is its unique maximizer.
From the E-step, we have the posterior probability $w_{i}(\eta)=\frac{p_{i}^{\star}\eta}{d_{i}(\eta)}$, which gives the identity:
$w_{i}(\eta)-\eta=\frac{\eta(1-\eta)(2p_{i}^{\star}-1)}{d_{i}(\eta)}.$
From this, we can express the derivative of the log-likelihood in terms of the EM update operator $T_{k}(\eta)$:
$\ell_{k}^{\prime}(\eta)=\sum_{i\in\mathcal{I}_{k}}\frac{2p_{i}^{\star}-1}{d_{i}(\eta)}=\frac{1}{\eta(1-\eta)}\sum_{i\in\mathcal{I}_{k}}\big(w_{i}(\eta)-\eta\big)=\frac{N_{k}}{\eta(1-\eta)}\big(T_{k}(\eta)-\eta\big).$
Therefore, the stationary point where $\ell_{k}^{\prime}(\eta)=0$ is equivalent to the fixed point where $T_{k}(\eta)=\eta$. This establishes that the unique maximizer $\widehat{\eta}_{k}$ is also the unique fixed point of $T_{k}$. The full-batch EM algorithm guarantees monotone ascent, i.e., $\ell_{k}(\eta_{k}^{(t+1)})\geq\ell_{k}(\eta_{k}^{(t)})$, and maps the interval $(0,1)$ to itself. Consequently, the sequence of estimates $\{\eta_{k}^{(t)}\}$ converges to the unique fixed point $\widehat{\eta}_{k}$. By Lemma 4.1, we know the true reliability $\eta_{k}^{\star}$ is a fixed point of $T_{k}$; uniqueness then implies that $\widehat{\eta}_{k}=\eta_{k}^{\star}$, so the EM iterates converge to the true reliability: $\eta_{k}^{(t)}\to\eta_{k}^{\star}$.
Remark E.1.
If $p_{i}^{\star}=\tfrac{1}{2}$ for all $i\in\mathcal{I}_{k}$, then $\ell_{k}(\eta)$ is constant on $(0,1)$, which implies $w_{i}(\eta)\equiv\eta$ and $T_{k}(\eta)\equiv\eta$. In this case, every $\eta\in(0,1)$ is a fixed point and uniqueness fails, as the model’s predictions are completely uninformative for assessing the reliability of annotator $k$.
Appendix F Resources
Models.
Datasets.